第4章離散確率変数
4.6 Poisson Distribution4.6 ポアソン分布
There are two main characteristics of a Poisson experiment.
ポアソン実験には二つの主な特徴がある。
- The Poisson probability distribution gives the probability of a number of events occurring in a fixed interval of time or space if these events happen with a known average rate and independently of the time since the last event. For example, a book editor might be interested in the number of words spelled incorrectly in a particular book. It might be that, on the average, there are five words spelled incorrectly in 100 pages. The interval is the 100 pages.
英語のヒント
ポアソン確率分布は、既知の平均発生率で、かつ前回の事象からの経過時間と独立に事象が起こる場合に、一定の時間または空間の区間で事象が何回起こるかの確率を与える。たとえば、書籍の編集者は、ある本で綴りを誤った単語の数に関心をもつかもしれない。平均して100ページに5語の綴り間違いがあるとする。この場合、区間は100ページである。 - The Poisson distribution may be used to approximate the binomial if the probability of success is "small" (such as 0.01) and the number of trials is "large" (such as 1,000). You will verify the relationship in the homework exercises. n is the number of trials, and p is the probability of a "success."
英語のヒント
成功確率が「小さく」(たとえば0.01)、試行回数が「大きい」(たとえば1,000)とき、ポアソン分布で二項分布を近似できる。この関係は宿題で確認する。nは試行回数、pは「成功」の確率である。
The random variable X = the number of occurrences in the interval of interest.
確率変数X = 関心のある区間内での発生回数である。
Example 4.26例題 4.26
The average number of loaves of bread put on a shelf in a bakery in a half-hour period is 12. Of interest is the number of loaves of bread put on the shelf in five minutes. The time interval of interest is five minutes. What is the probability that the number of loaves, selected randomly, put on the shelf in five minutes is three?
英語のヒント
パン屋の棚に30分間で並べられるパンは、平均12個である。5分間で棚に並べられるパンの個数に関心がある。関心のある時間区間は5分である。無作為に選んだ5分間に、棚に並べられるパンが3個である確率はいくらか。
Let X = the number of loaves of bread put on the shelf in five minutes. If the average number of loaves put on the shelf in 30 minutes (half-hour) is 12, then the average number of loaves put on the shelf in five minutes is (12) = 2 loaves of bread.
X = 5分間で棚に並べられるパンの個数とする。30分(半時間)で並べられる平均個数が12なら、5分間の平均個数は(12) = 2個である。
The probability question asks you to find P(x = 3).
この確率の問いではP(x = 3)を求める。
Example 4.27例題 4.27
A bank expects to receive six bad checks per day, on average. What is the probability of the bank getting fewer than five bad checks on any given day? Of interest is the number of checks the bank receives in one day, so the time interval of interest is one day. Let X = the number of bad checks the bank receives in one day. If the bank expects to receive six bad checks per day then the average is six checks per day. Write a mathematical statement for the probability question.
英語のヒント
ある銀行では、1日当たり平均6枚の不渡り小切手を受け取ると予想される。ある日の不渡り小切手が5枚未満である確率はいくらか。関心があるのは1日に受け取る小切手の枚数なので、時間区間は1日である。X = 銀行が1日に受け取る不渡り小切手の枚数とする。1日6枚と予想されるので、平均は1日6枚である。この確率の問いを数学的に表せ。
Solution解き方
P(x < 5)
Example 4.28例題 4.28
You notice that a news reporter says "uh," on average, two times per broadcast. What is the probability that the news reporter says "uh" more than two times per broadcast.
英語のヒント
あるニュース記者が、放送1回につき平均2回「えー」と言うことに気づいた。1回の放送で「えー」と2回より多く言う確率はいくらか。
This is a Poisson problem because you are interested in knowing the number of times the news reporter says "uh" during a broadcast.
1回の放送中に記者が「えー」と言う回数に関心があるので、これはポアソン分布の問題である。
a. What is the interval of interest?
a. 関心のある区間は何か。
b. What is the average number of times the news reporter says "uh" during one broadcast?
b. 1回の放送中に記者が「えー」と言う平均回数はいくらか。
c. Let X = ____________. What values does X take on?
c. X = ____________とする。Xはどのような値を取るか。
d. The probability question is P(______).
d. 求める確率はP(______)である。
Solution解き方
a. one broadcast
a. 1回の放送
b. 2
c. Let X = the number of times the news reporter says "uh" during one broadcast.
x = 0, 1, 2, 3, ...
x = 0, 1, 2, 3, ...
c. X = 1回の放送中に記者が「えー」と言う回数とする。x = 0, 1, 2, 3, …。
d. P(x > 2)
Notation for the Poisson: P = Poisson Probability Distribution Functionポアソン分布の表記:P = ポアソン確率分布関数
X ~ P(μ)
Read this as "X is a random variable with a Poisson distribution." The parameter is μ (or λ); μ (or λ) = the mean for the interval of interest. The standard deviation of the Poisson distribution with mean µ is σ=√μ
これを「Xはポアソン分布に従う確率変数である」と読む。母数はμ(またはλ)で、μ(またはλ)= 関心のある区間での平均である。平均μのポアソン分布の標準偏差はσ = √μである。
Example 4.29例題 4.29
Leah receives about six telephone calls between 8 a.m. and 10 a.m. What is the probability that Leah receives more than one call in the next 15 minutes?
リアには午前8時から10時までに約6件の電話がかかってくる。次の15分間に1件より多く電話を受ける確率はいくらか。
Let X = the number of calls Leah receives in 15 minutes. (The interval of interest is 15 minutes or hour.)
X = リアが15分間に受ける電話の件数とする(関心のある区間は15分、すなわち時間)。
x = 0, 1, 2, 3, ...
If Leah receives, on the average, six telephone calls in two hours, and there are eight 15 minute intervals in two hours, then Leah receives
リアが2時間で平均6件の電話を受け、2時間には15分の区間が八つあるので、15分間に受ける電話は、
(6) = 0.75 calls in 15 minutes, on average. So, μ = 0.75 for this problem.
平均(6) = 0.75件である。したがって、この問題ではμ = 0.75となる。
Using the Calculator電卓の操作
The Poisson distribution is discrete, and thus > 1 includes all whole numbers through infinity. The solution is to subtract the probability less than 1 thus
ポアソン分布は離散型なので、> 1は無限までのすべての整数を含む。1未満の確率を差し引くことで解き、とする。
Figure 4.7図 4.7
The y-axis in Figure 4.7 contains the probability of x where X = the number of calls in 15 minutes.
図4.7のy軸はxの確率を表す。ここでX = 15分間の電話の件数である。
Try Itやってみよう
A customer service center receives about ten emails every half-hour. What is the probability that the customer service center receives more than four emails in the next six minutes?
顧客サービスセンターには、30分ごとに約10通のメールが届く。次の6分間に4通より多く届く確率はいくらか。
Example 4.30例題 4.30
According to Baydin, an email management company, an email user gets, on average, 147 emails per day. Let X = the number of emails an email user receives per day. The discrete random variable X takes on the values x = 0, 1, 2 …. The random variable X has a Poisson distribution: X ~ P(147). The mean is 147 emails.
メール管理会社Baydinによると、メール利用者は1日平均147通のメールを受け取る。X = 利用者が1日に受け取るメールの通数とする。離散確率変数Xはx = 0, 1, 2, …の値を取る。Xはポアソン分布に従い、X ~ P(147)である。平均は147通である。
- What is the probability that an email user receives exactly 160 emails per day?メール利用者が1日にちょうど160通受け取る確率はいくらか。
- What is the probability that an email user receives at most 160 emails per day?メール利用者が1日に受け取るメールが160通以下である確率はいくらか。
- What is the standard deviation?標準偏差はいくらか。
Solution解き方
- P(x = 160) = poissonpdf(147, 160) ≈ 0.0180
- P(x ≤ 160) = poissoncdf(147, 160) ≈ 0.8666
- Standard Deviation =標準偏差 =
Try Itやってみよう
According to a recent poll by the Pew Internet Project, people between the ages of 14 and 17 send an average of 187 text messages each day. Let X = the number of texts that a girl aged 14 to 17 sends per day. The discrete random variable X takes on the values x = 0, 1, 2 …. The random variable X has a Poisson distribution: X ~ P(187). The mean is 187 text messages.
ピュー・インターネット・プロジェクトの最近の調査によると、14〜17歳の人は1日平均187通のテキストメッセージを送る。X = 14〜17歳の少女が1日に送るメッセージの通数とする。離散確率変数Xはx = 0, 1, 2, …の値を取る。Xはポアソン分布に従い、X ~ P(187)である。平均は187通である。
- What is the probability that a person sends exactly 175 texts per day?ある人が1日にちょうど175通のメッセージを送る確率はいくらか。
- What is the probability that a person sends at most 150 texts per day?ある人が1日に送るメッセージが150通以下である確率はいくらか。
- What is the standard deviation?標準偏差はいくらか。
Example 4.31例題 4.31
Text message users receive or send an average of 41.5 text messages per day.
テキストメッセージ利用者は、1日平均41.5通を送信または受信する。
- How many text messages does a text message user receive or send per hour?利用者は1時間当たり何通のテキストメッセージを受信または送信するか。
- What is the probability that a text message user receives or sends two messages per hour?利用者が1時間に2通のメッセージを受信または送信する確率はいくらか。
- What is the probability that a text message user receives or sends more than two messages per hour?利用者が1時間に2通より多くのメッセージを受信または送信する確率はいくらか。
Solution解き方
- Let X = the number of texts that a user sends or receives in one hour. The average number of texts received per hour is ≈ 1.7292.X = 利用者が1時間に送信または受信するメッセージの通数とする。1時間当たりの平均受信数は ≈ 1.7292である。
- X ~ P(1.7292), so P(x = 2) = poissonpdf(1.7292, 2) ≈ 0.2653X ~ P(1.7292)なので、P(x = 2) = poissonpdf(1.7292, 2) ≈ 0.2653となる。
- P(x > 2) = 1 – P(x ≤ 2) = 1 – poissoncdf(1.7292, 2) ≈ 1 – 0.7495 = 0.2505P(x > 2) = 1 − P(x ≤ 2) = 1 − poissoncdf(1.7292, 2) ≈ 1 − 0.7495 = 0.2505
Try Itやってみよう
Atlanta’s Hartsfield-Jackson International Airport is the busiest airport in the world. On average there are 2,700 arrivals and departures each day.
アトランタのハーツフィールド・ジャクソン国際空港は、世界で最も利用の多い空港である。1日平均2,700回の到着と出発がある。
- How many airplanes arrive and depart the airport per hour?1時間当たり何機の飛行機が到着・出発するか。
- What is the probability that there are exactly 100 arrivals and departures in one hour?1時間にちょうど100回の到着・出発がある確率はいくらか。
- What is the probability that there are at most 100 arrivals and departures in one hour?1時間の到着・出発が100回以下である確率はいくらか。
Example 4.32例題 4.32
On a certain day in May starting at 4:30 PM, the probability of low seismic activity for the next 48 hours in Alaska was reported as about 1.02%. Use this information for the next 200 days to find the probability that there will be low seismic activity in ten of the next 200 days. Use both the binomial and Poisson distributions to calculate the probabilities. Are they close?
5月のある日、午後4時30分からの48時間以内にアラスカで小規模な地震活動が起こる確率は約1.02%と報告された。この情報を今後200日間に用い、そのうち10日で小規模な地震活動がある確率を求めよ。二項分布とポアソン分布の両方で計算せよ。値は近いか。
Solution解き方
Let X = the number of days with low seismic activity.
X = 小規模な地震活動がある日数とする。
Using the binomial distribution:
二項分布を使うと、
- P(x = 10) = binompdf(200, .0102, 10) ≈ 0.000039
Using the Poisson distribution:
ポアソン分布を使うと、
- Calculate μ = np = 200(0.0102) ≈ 2.04μ = np = 200(0.0102) ≈ 2.04を計算する。
- P(x = 10) = poissonpdf(2.04, 10) ≈ 0.000045
We expect the approximation to be good because n is large (greater than 20) and p is small (less than 0.05). The results are close—both probabilities reported are almost 0.
nが大きく(20より大きい)、pが小さい(0.05未満)ため、よい近似になると予想される。結果は近く、どちらの確率もほぼ0である。
Try Itやってみよう
On a certain day in May starting at 4:30 PM, the probability of moderate seismic activity for the next 48 hours in the Kuril Islands off the coast of Japan was reported at about 1.43%. Use this information for the next 100 days to find the probability that there will be low seismic activity in five of the next 100 days. Use both the binomial and Poisson distributions to calculate the probabilities. Are they close?
5月のある日、午後4時30分からの48時間以内に、日本沖の千島列島で中程度の地震活動が起こる確率は約1.43%と報告された。この情報を今後100日間に用い、そのうち5日で小規模な地震活動がある確率を求めよ。二項分布とポアソン分布の両方で計算せよ。値は近いか。
